# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 3.2.4 Mamba-based RSISR
Recently, empowered by state space model (SSM) involving near-linear complexity modeling long-distance dependencies, Mamba [243] has demonstrated superior performance compared to Transformer in natural language processing (NLP). Specifically, Mamba introduces a selective mechanism into the traditional SSM to improve computational efficiency and significantly enhance the ability to model long sequences. Building on success in sequence modeling, Mamba has recently demonstrated remarkable potential across diverse vision domains, achieving state-of-the-art performance in fundamental tasks, including image classification [244], object detection [245], and semantic segmentation [246]. However, the research on Mamba for low-level vision tasks is still in its infancy. In particular, a small number of studies have begun to use Mamba to perform RSISR tasks.
@@
In recent work, Zhu et al. [104] proposed a ConvMambaSRF method that integrates SSM and CNN for RSISR. Specifically, they used SSM and CNN to model global dependencies and extract local details respectively, and realized multi-level feature fusion through the global-detail reconstruction module (GDRM). Besides, Zhi et al. [247] proposed a MambaFormerSR method combining Mamba and Transformer for RSISR. They designed a state-space-attention fusion module (SAFM) and combined it with a convolutional Fourier feedforward network (CTFFN) to strengthen frequency domain attention. The spatial-frequency dual-path framework proposed by Xiao et al. [105] for RSISR uses a frequency selection module (FSM) and a state space module (VSSM) to extract high-frequency details and capture long-range spatial correlations, respectively, and further designs a hybrid gating module (HGM) to dynamically fuse multi-level features, achieving outstanding performance. To improve computational efficiency, Zhang and Wang [248] combined SSM with residual connections to implement an equivalent attention mechanism with linear complexity, and experimentally demonstrated that it significantly enhances cross-region feature associations while preserving local textures.
@@
Mamba-based RSISR method successfully solves the local receptive field limitations of traditional CNN and the high computational bottleneck of Transformer by leveraging the linear complexity global modeling capability of SSM, demonstrating a new paradigm for large-scale remote sensing image reconstruction. Although some studies, such as hybrid architecture design strategies [104, 247] and multi-modal feature collaboration [248], have significantly improved reconstruction performance, existing methods still face challenges such as blurred high-frequency details at extremely low-resolutions and insufficient temporal modeling of multi-phase images. In the future, the research on RSISR based on Mamba may focus on the following aspects: (a) lightweight 3D Mamba architecture design for time-series video RSISR and dynamic scene reconstruction, (b) physical degradation-aware SSM optimization, combined with imaging models to enhance the robustness of unknown degradation scenes, (c) cross-modal state space modeling, integrating multi-spectral, SAR and other multi-source remote sensing data to build a unified representation framework for heterogeneous features. With further breakthroughs in hardware perception algorithms and adaptive selectivity mechanisms, Mamba has the potential to become a universal model for RSISR tasks.
@@ 3.2.5 Diffusion Model-based RSISR
To the best of our knowledge, Liu et al. [249] proposed the first diffusion model-based RSISR method, which uses LR images as conditional information to generate HR images. Wu et al. [84] proposed an HSR-Diff method for RSISR, which uses conditional Transformer to fuse HR images and LR images, and then optimizes the reconstruction through hierarchical feature iteration. The efficient hybrid conditional diffusion model (EHC-DMSR) for RSISR proposed by Han et al. [250] combines Transformer-CNN to extract conditional features and introduces Fourier high frequency constraints to accelerate reasoning. In addition, Meng et al. [94] proposed the FastDiffSR method to improve the sampling strategy and reduce the number of diffusion steps. An et al.[95] proposed a lightweight diffusion model LWTDM, which uses a cross-attention encoder-decoder to simplify the denoising network and integrates denoising diffusion implicit models (DDIMs) to accelerate sampling. The DiffALS method [251] introduces a noise discriminator (ND) through an adversarial learning strategy to enhance the diversity of detail generation. Furthermore, the TCDM method proposed by Zhang et al. [252] designs a conditional truncated noise generator (CTNG) for large-scale super-resolution, and embeds pixel-level constraints in combination with the texture consistency diffusion process to accelerate reasoning. Xiao et al. [98] proposed an EDiffSR method for RSISR, which improves efficiency by simplifying channel attention. Some methods focus on semantics and global optimization. For example, Wang et al. [111] combined vector map semantic guidance and proposed a Semantic Guided Diffusion Model (SGDM). Zhu et al. [253] introduced low-rank adaptation (LoRA) to alleviate the distribution difference between natural images and RSI. Weng et al. [110] proposed an image reconstruction representation-diffusion model for RSISR, which uses a pre-trained encoder to guide high-frequency reconstruction.
@@ 3.2.6 Hybrid RSISR Methods
In recent years, some studies have adopted hybrid architectures to improve performance of RSISR. For example, He et al. [254] proposed a hybird model (DsTer) that combines densely connected Transformer and ResNet, which enhances spectral super-resolution capabilities through multi-level Transformer feature fusion and demonstrates its advantages in long-range dependency modeling on natural and remote sensing datasets. In addition, Tu et al. [255] proposed a GAN method (SWCGAN) combining Swin Transformer and CNN for RSISR, which improves the multi-scale feature extraction capability through residual dense blocks and outperforms the traditional CNN method on the UCMerced dataset. However, the adversarial training of the SWCGAN method is prone to introduce artifacts. Wang et al. [256] proposed a dilated Transformer generative adversarial network (DTGAN), which introduced dilated convolution into the self-attention module of the Transformer and dynamically focused on local and global features of different scales, significantly improving the performance of land classification tasks.
@@
Recently, some studies have focused on exploring the generation ability and stability of hybrid architectures. For example, Guo et al. [262] proposed a sparse-activated sub-pixel transformer network (SSTNet), which enhances edge sharpness and texture details through sparse activation in sub-pixel space, but a sparse constraint term requires additional design. Besides, Liu et al. [263] proposed an improved RSISR method, which designs a multi-stage hybrid Transformer generator and combines Charbonnier loss with TV loss to improve training stability. The Fusing Transformers and CNN method (ConvFormerSR) for RSISR proposed by Li et al. [264] designed a cross-sensor feature fusion module (FFM) and spectral loss function, and achieved high-spectral consistency reconstruction of Landsat-8 and Sentinel-2 data by enhancing the high-order spatial interaction of Transformer. MSWAGAN method [265] designs a multi-scale sliding window attention (MSWA) to capture local multi-scale features, and combines the Transformer to model long-range pixel associations without increasing the number of parameters. Zhu et al. [108] proposed a Transformer-based Multi-modal Generative Adversarial Network (TMGAN) for multimodal data, which designed a global Transformer generator and fused depth and spectral information through self-attention to improve performance. The Enhanced Swin Transformer with U-Net GAN (ESTUGAN) proposed by Yu et al. [266] employs a Swin Transformer-based generator and a U-Net discriminator, which suppresses artifacts through region-aware adversarial learning. Specifically, ESTUGAN uses Best-buddy loss and Back-projection loss to enhance reconstruction fidelity. In the Trans-CNN GAN method proposed by Lin et al. [267], the global attention of the Transformer is used to compensate for the local defects of the generator. Huo et al. [268] proposed a STGAN model based on GANs and self-attention mechanism, which achieved better high-frequency detail recovery ability in RSISR through multi-CNN-Swin Transformer module (MCST) and improved association attention module (RAM-V).
@@ 4 Unsupervised RSISR Methods
Supervised RSISR methods usually need massive amounts of labeled training data, which face many challenges in practical applications, such as high cost of acquiring real HR data, poor annotation transferability across sensor scenarios, and difficulty in eliminating deviations between simulated data and real physical degradation models. In contrast, unsupervised methods alleviate the strict reliance on labeled data, which leverage the intrinsic characteristics of images, such as self-similarity, multi-source information complementarity, and domain-specific physical priors. Therefore, it is still necessary to conduct research on unsupervised RSISR methods. In this section, we conduct a detailed investigation on the unsupervised RSISR methods from the perspectives of traditional methods and deep learning.
@@ 4.1 Traditional RSISR Methods
Although deep learning has emerged as the dominant approach for RSISR, traditional unsupervised methods still hold significant research value due to their demonstrated advantages such as model robustness, physical interpretability, and cross-scenario generalization capabilities. In addition, these conventional techniques remain particularly promising in resource-constrained applications, particularly in scenarios involving small-sample learning, low-resource consumption, and specialized tasks like zero-shot super-resolution. Furthermore, their well-established theoretical frameworks provide offer valuable insights for advancing mainstream algorithms through improved architecture design and interpretability enahncement. Generally, traditional unsupervised RSISR mainly focuses on reconstruction-based methods, which involve the following aspects: interpolation-based, frequency domain-based, iterative back projection (IBP)-based, and regularization constraint-based.
@@ 4.1.1 Interpolation-based Methods
In early work [269], wavelet transform and spatial interpolation were combined to improve reconstruction performance, which used a spectrum separation processing strategy to improve resolution while suppressing interpolation artifacts. In additon, Zhou et al. [44] proposed an interpolation super-resolution method based on multi-surface fitting, which focuses on spatial structure-driven surface modeling and maximum a posteriori probability fusion to achieve high-frequency detail preservation. To address the problems of blur and block effects caused by the loss of high-frequency information in conventional interpolation, Han et al. [270] proposed a RSISR framework that combines frequency domain segmentation enhancement with spatial domain interpolation optimization based on bicubic interpolation and discrete wavelet transform (DWT), which significantly improves the reconstruction performance and preserves the details of the ground objects. A similar work is the multi-band collaborative super-resolution framework based on dual-tree complex wavelet transform (DT-CWT) and improved edge-directed interpolation (INEDI) proposed by Solanki et al. [55]. In addition, Mareboyana and Moigne [271] proposed an edge-directed radial basis function (EDRBF) driven super-resolution method, which injects edge-constrained radial basis adaptive interpolation mechanism and co-optimizes sub-pixel registration, showing strong robustness to registration errors.
@@ 4.1.2 POCS-based Methods
Projection onto Convex Sets (POCS) is an iterative optimization method based on convex set constraints, which approximates the feasible solution of RSISR reconstruction by alternately projecting onto a priori knowledge constraint sets of the degradation model. Earlier works [272, 273] have demonstrated that POCS has significant potential in improving super-resolution performance. A more concrete work is the two-stage regularization framework based on POCS proposed by Aguena and Mascarenhas [274] for multispectral-panchromatic fusion, which achieves spectral-spatial feature collaborative enhancement through sequential and parallel projection or least squares synthesis. In addition, Xie et al. [275] proposed a robust POCS iterative framework based on improved sub-pixel shift estimation for blind image super-resolution in 2009, but it has the disadvantage of slow convergence speed. In order to solve the problem of imbalanced image data of the Chang’e-1 three-line array, Zhang et al. [43] proposed a RSISR method based on global weighted POCS, which significantly improved the performance of lunar surface topography reconstruction. In 2016, Liu et al. [52] proposed an improved POCS algorithm based on visual mechanism for RSISR, which achieves infrared image target edge enhancement and background noise suppression through variable correction threshold consistency constraint and human visual contrast constraint. In order to solve the problem of inaccurate PSF estimation in traditional POCS, Fan et al. [276] proposed a slant-edge estimation strategy based on the linear relationship of PSF of LR images, which greatly improved the performance of image reconstruction. In addition, Dai et al. [277] proposed an improved POCS algorithm for RSISR, which employs iterative curvature-based interpolation (ICBI) to generate high-resolution initial estimates, effectively alleviating the edge blur and detail loss problems in POCS reconstruction. A similar work is that Wang et al. [75] proposed to use gradient interpolation instead of nearest neighbor interpolation to solve the problem of edge blur and iterative subjectivity in hyperspectral super-resolution, and designed an adaptive stopping criterion based on the mean square error of adjacent iterations to improve the quality of full-band reconstruction.
@@ 4.1.3 IBP-based Methods
Iterative Back Projection (IBP)-based super-resolution methods mainly improve the resolution of remote sensing images through error feedback iterative optimization . To the best of our knowledge, IBP was first introduced as a multi-frame super-resolution algorithm in the work of Irani and Peleg et al. [278]. Early work in [279] proposed a sub-pixel migration multi-observation model, which established super-resolution equations and employed the IBP algorithm to obtain HR images. Besides, Li et al. [280] proposed an improved IBP algorithm for RSISR by integrating inverse-forward elastic registration with block-wise processing to address local affine transformations, which enhanced geometric adaptability and high-frequency preservation. Their subsequent work [281] further demonstrated the effectiveness in stereo image fusion of ALOS satellite’s PRISM sensors. A similar work is that Yan and Lu [282] combined IBP with the Papoulis-Gerchberg extrapolation method to solve the anisotropic resolution problem of MRI.

### v3

@@ 3.2.4 Mamba-based RSISR
Recently, empowered by a state space model (SSM) involving near-linear complexity modeling long-distance dependencies, Mamba [246] has demonstrated superior performance compared to Transformer in natural language processing (NLP). Specifically, Mamba introduces a selective mechanism into the traditional SSM to improve computational efficiency and significantly enhance the ability to model long sequences. Building on success in sequence modeling, Mamba has recently demonstrated remarkable potential across diverse vision domains, achieving state-of-the-art performance in fundamental tasks, including image classification [247], object detection [248], and semantic segmentation [249]. However, the research on Mamba for low-level vision tasks is still in its infancy. In particular, a small number of studies have begun to use Mamba to perform RSISR tasks.
@@
In recent work, Zhu et al. [105] proposed a ConvMambaSRF method that integrates SSM and CNN for RSISR. Specifically, they used SSM and CNN to model global dependencies and extract local details, respectively, and realized multi-level feature fusion through the global-detail reconstruction module (GDRM). Besides, Zhi et al. [250] proposed a MambaFormerSR method combining Mamba and Transformer for RSISR. They designed a state-space-attention fusion module (SAFM) and combined it with a convolutional Fourier feedforward network (CTFFN) to strengthen frequency domain attention. The spatial-frequency dual-path framework proposed by Xiao et al. [106] for RSISR uses a frequency selection module (FSM) and a state space module (VSSM) to extract high-frequency details and capture long-range spatial correlations, respectively, and further designs a hybrid gating module (HGM) to dynamically fuse multi-level features, achieving outstanding performance. To improve computational efficiency, Zhang and Wang [251] combined SSM with residual connections to implement an equivalent attention mechanism with linear complexity, and experimentally demonstrated that it significantly enhances cross-region feature associations while preserving local textures.
@@
Mamba-based RSISR method successfully solves the local receptive field limitations of traditional CNN and the high computational bottleneck of Transformer by leveraging the linear complexity global modeling capability of SSM, demonstrating a new paradigm for large-scale RSI reconstruction. Although some studies, such as hybrid architecture design strategies [105, 250] and multi-modal feature collaboration [251], have significantly improved reconstruction performance, existing methods still face challenges such as blurred high-frequency details at extremely low-resolution and insufficient temporal modeling of multi-phase images. In the future, the research on RSISR based on Mamba may focus on the following aspects: (a) lightweight 3D Mamba architecture design for time-series video RSISR and dynamic scene reconstruction, (b) physical degradation-aware SSM optimization, combined with imaging models to enhance the robustness of unknown degradation scenes, (c) cross-modal state space modeling, integrating multi-spectral, SAR, and other multi-source remote sensing data to build a unified representation framework for heterogeneous features. With further breakthroughs in hardware perception algorithms and adaptive selectivity mechanisms, Mamba has the potential to become a universal model for RSISR tasks.
@@ 3.2.5 Diffusion Model-based RSISR
To the best of our knowledge, Liu et al. [252] proposed the first diffusion model-based RSISR method, which uses LR images as conditional information to generate HR images. Wu et al. [85] proposed an HSR-Diff method for RSISR, which uses a conditional Transformer to fuse HR images and LR images, and then optimizes the reconstruction through hierarchical feature iteration. The efficient hybrid conditional diffusion model (EHC-DMSR) for RSISR proposed by Han et al. [253] combines Transformer-CNN to extract conditional features and introduces Fourier high-frequency constraints to accelerate reasoning. In addition, Meng et al. [95] proposed the FastDiffSR method to improve the sampling strategy and reduce the number of diffusion steps. An et al.[96] proposed a lightweight diffusion model LWTDM, which uses a cross-attention encoder-decoder to simplify the denoising network and integrates denoising diffusion implicit models (DDIMs) to accelerate sampling. The DiffALS method [254] introduces a noise discriminator (ND) through an adversarial learning strategy to enhance the diversity of detail generation. Furthermore, the TCDM method proposed by Zhang et al. [255] designs a conditional truncated noise generator (CTNG) for large-scale super-resolution and embeds pixel-level constraints in combination with the texture consistency diffusion process to accelerate reasoning. Xiao et al. [99] proposed an EDiffSR method for RSISR, which improves efficiency by simplifying channel attention. Some methods focus on semantics and global optimization. For example, Wang et al. [112] combined vector map semantic guidance and proposed a Semantic Guided Diffusion Model (SGDM). Zhu et al. [256] introduced low-rank adaptation (LoRA) to alleviate the distribution difference between natural images and RSI. Weng et al. [111] proposed an image reconstruction representation-diffusion model for RSISR, which uses a pre-trained encoder to guide high-frequency reconstruction.
@@ 3.2.6 Hybrid RSISR Methods
In recent years, some studies have adopted hybrid architectures to improve the performance of RSISR. For example, He et al. [257] proposed a hybrid model (DsTer) that combines densely connected Transformer and ResNet, which enhances spectral super-resolution capabilities through multi-level Transformer feature fusion and demonstrates its advantages in long-range dependency modeling on natural and remote sensing datasets. In addition, Tu et al. [258] proposed a GAN method (SWCGAN) combining Swin Transformer and CNN for RSISR, which improves the multi-scale feature extraction capability through residual dense blocks and outperforms the traditional CNN method on the UCMerced dataset. However, the adversarial training of the SWCGAN method is prone to introducing artifacts. Wang et al. [259] proposed a dilated Transformer generative adversarial network (DTGAN), which introduced dilated convolution into the self-attention module of the Transformer and dynamically focused on local and global features of different scales, significantly improving the performance of land classification tasks.
@@
Recently, some studies have focused on exploring the generation ability and stability of hybrid architectures. For example, Guo et al. [265] proposed a sparse-activated sub-pixel transformer network (SSTNet), which enhances edge sharpness and texture details through sparse activation in sub-pixel space, but a sparse constraint term requires additional design. Besides, Liu et al. [266] proposed an improved RSISR method, which designs a multi-stage hybrid Transformer generator and combines Charbonnier loss with TV loss to improve training stability. The Fusing Transformers and CNN method (ConvFormerSR) for RSISR proposed by Li et al. [267] designed a cross-sensor feature fusion module (FFM) and spectral loss function, and achieved high-spectral consistency reconstruction of Landsat-8 and Sentinel-2 data by enhancing the high-order spatial interaction of Transformer. MSWAGAN method [268] designs a multi-scale sliding window attention (MSWA) to capture local multi-scale features, and combines the Transformer to model long-range pixel associations without increasing the number of parameters. Zhu et al. [109] proposed a Transformer-based Multi-modal Generative Adversarial Network (TMGAN) for multimodal data, which designed a global Transformer generator and fused depth and spectral information through self-attention to improve performance. The Enhanced Swin Transformer with U-Net GAN (ESTUGAN) proposed by Yu et al. [269] employs a Swin Transformer-based generator and a U-Net discriminator, which suppresses artifacts through region-aware adversarial learning. Specifically, ESTUGAN uses Best-buddy loss and Back-projection loss to enhance reconstruction fidelity. In the Trans-CNN GAN method proposed by Lin et al. [270], the global attention of the Transformer is used to compensate for the local defects of the generator. Huo et al. [271] proposed a STGAN model based on GANs and self-attention mechanism, which achieved better high-frequency detail recovery ability in RSISR through a multi-CNN-Swin Transformer module (MCST) and an improved association attention module (RAM-V).
@@ 4 Unsupervised RSISR Methods
Supervised RSISR methods usually need massive amounts of labeled training data, which face many challenges in practical applications, such as high cost of acquiring real HR data, poor annotation transferability across sensor scenarios, and difficulty in eliminating deviations between simulated data and real physical degradation models. In contrast, unsupervised methods alleviate the strict reliance on labeled data, which leverage the intrinsic characteristics of images, such as self-similarity, multi-source information complementarity, and domain-specific physical priors. Therefore, it is still necessary to conduct research on unsupervised RSISR methods. In this section, we conduct a detailed investigation of the unsupervised RSISR methods from the perspectives of traditional methods and deep learning.
@@ 4.1 Traditional RSISR Methods
Although deep learning has emerged as the dominant approach for RSISR, traditional unsupervised methods still hold significant research value due to their demonstrated advantages, such as model robustness, physical interpretability, and cross-scenario generalization capabilities. In addition, these conventional techniques remain particularly promising in resource-constrained applications, particularly in scenarios involving small-sample learning, low-resource consumption, and specialized tasks like zero-shot super-resolution. Furthermore, their well-established theoretical frameworks offer valuable insights for advancing mainstream algorithms through improved architecture design and interpretability enhancement. Generally, traditional unsupervised RSISR mainly focuses on reconstruction-based methods, which involve the following aspects: interpolation-based, frequency domain-based, iterative back projection (IBP)-based, and regularization constraint-based.
@@ 4.1.1 Interpolation-based Methods
In early work [272], wavelet transform and spatial interpolation were combined to improve reconstruction performance, which used a spectrum separation processing strategy to improve resolution while suppressing interpolation artifacts. In addition, Zhou et al. [45] proposed an interpolation super-resolution method based on multi-surface fitting, which focuses on spatial structure-driven surface modeling and maximum a posteriori probability fusion to achieve high-frequency detail preservation. To address the problems of blur and block effects caused by the loss of high-frequency information in conventional interpolation, Han et al. [273] proposed a RSISR framework that combines frequency domain segmentation enhancement with spatial domain interpolation optimization based on bicubic interpolation and discrete wavelet transform (DWT), which significantly improves the reconstruction performance and preserves the details of the ground objects. A similar work is the multi-band collaborative super-resolution framework based on dual-tree complex wavelet transform (DT-CWT) and improved edge-directed interpolation (INEDI) proposed by Solanki et al. [56]. In addition, Mareboyana and Moigne [274] proposed an edge-directed radial basis function (EDRBF) driven super-resolution method, which injects an edge-constrained radial basis adaptive interpolation mechanism and co-optimizes sub-pixel registration, showing strong robustness to registration errors.
@@ 4.1.2 POCS-based Methods
Projection onto Convex Sets (POCS) is an iterative optimization method based on convex set constraints, which approximates the feasible solution of RSISR reconstruction by alternately projecting onto a priori knowledge constraint sets of the degradation model. Earlier works [275, 276] have demonstrated that POCS has significant potential in improving super-resolution performance. A more concrete work is the two-stage regularization framework based on POCS proposed by Aguena and Mascarenhas [277] for multispectral-panchromatic fusion, which achieves spectral-spatial feature collaborative enhancement through sequential and parallel projection or least squares synthesis. In addition, Xie et al. [278] proposed a robust POCS iterative framework based on improved sub-pixel shift estimation for blind image super-resolution in 2009, but it has the disadvantage of slow convergence speed. In order to solve the problem of imbalanced image data of the Chang’e-1 three-line array, Zhang et al. [44] proposed an RSISR method based on global weighted POCS, which significantly improved the performance of lunar surface topography reconstruction. In 2016, Liu et al. [53] proposed an improved POCS algorithm based on visual mechanism for RSISR, which achieves infrared image target edge enhancement and background noise suppression through variable correction threshold consistency constraint and human visual contrast constraint. In order to solve the problem of inaccurate PSF estimation in traditional POCS, Fan et al. [279] proposed a slant-edge estimation strategy based on the linear relationship of PSF of LR images, which greatly improved the performance of image reconstruction. In addition, Dai et al. [280] proposed an improved POCS algorithm for RSISR, which employs iterative curvature-based interpolation (ICBI) to generate high-resolution initial estimates, effectively alleviating the edge blur and detail loss problems in POCS reconstruction. A similar work is that Wang et al. [76] proposed to use gradient interpolation instead of nearest neighbor interpolation to solve the problem of edge blur and iterative subjectivity in hyperspectral super-resolution, and designed an adaptive stopping criterion based on the mean square error of adjacent iterations to improve the quality of full-band reconstruction.
@@ 4.1.3 IBP-based Methods
Iterative Back Projection (IBP)-based super-resolution methods mainly improve the resolution of RSIs through error feedback iterative optimization. To the best of our knowledge, IBP was first introduced as a multi-frame super-resolution algorithm in the work of Irani and Peleg et al. [281]. Early work in [282] proposed a sub-pixel migration multi-observation model, which established super-resolution equations and employed the IBP algorithm to obtain HR images. Besides, Li et al. [283] proposed an improved IBP algorithm for RSISR by integrating inverse-forward elastic registration with block-wise processing to address local affine transformations, which enhanced geometric adaptability and high-frequency preservation. Their subsequent work [284] further demonstrated the effectiveness in stereo image fusion of ALOS satellite’s PRISM sensors. A similar work is that Yan and Lu [285] combined IBP with the Papoulis-Gerchberg extrapolation method to solve the anisotropic resolution problem of MRI.